OpenStudy (anonymous):

COUNTRY A FAMILY NET WORTH WAS 340,00 AND THERE WHERE 8.8 X 10^7 FAMILIES CALCUALTE FAMILY NETWORL IN COMMUNITY A USE SCIENTIFIC NOTATION

Directrix (directrix):

340,00 AND THERE WHERE 8.8 X 10^7 FAMILIES. 3.4 * 10^2 * 8.8 * 10^7 = 3.4 * 8.8 * 10^2 * 10^7 = 29.92 * 10^9 = 2.992* 10^1 * 10 ^9 = 2.992 * 10^10 @termekq Please check for accuracy. Thanks.